OpenStudy (anonymous):

simplify, do not use negative exponents in your answer y^-5* y

OpenStudy (alexwee123):

y^-5* y =y^(-5+1) =y^-4 =1/y^4

OpenStudy (anonymous):

thank you! for some reason I had a brain fart when I did this question on my own and forgot that the five was negative. so my answer was 1/y^6. which obviously was wrong. thanks :)